OpenStudy (anonymous):

Part 1: Change 4x – 3y = –15 to slope-intercept form. You must show all work. Part 2: State the slope and y-intercept

OpenStudy (anonymous):

Slope intercept form. y=mx+b. Let y be on the right side and all other terms on the left. -3y = -15 -4x. Change signs by multiplying each term by -1 to make the y term positive. 3y = 4x + 15. Divide each term by 3, y = 4x/3 + 5. Final answer.